OpenStudy (anonymous):

An 11.2 g bullet leaves the muzzle of a rifle with a speed of 408.3 m/s. What constant force is exerted on the bullet while it is traveling down the 0.6 m length of the barrel of the rifle? Answer in units of N

OpenStudy (jadzia):

for the first one you'll need to know F=ma to solve for the constant force but take note, you are given the speed not acceleration so you still have to solve for acceleration Use the ffg formula \(V_f^2 - V_o^2 = 2as\) where Vf = velocity of the bullet as it leaves the muzzle = 408.3 m/sec (given) Vo = initial velocity of bullet = 0 a = acceleration of bullet s = length of the rifle's barrel =0.6 m after calculating for acceleration, substitute the value into the formual F=ma and solve for the force

OpenStudy (anonymous):

Thank you for the help I tried all of this and got the wrong answer is there anyway you could write out the work with the answer so I could see what I did wrong?

OpenStudy (jadzia):

can you show me your work instead? I'll just check your solution if it's right

OpenStudy (jadzia):

given: m= 11.2 g Vf= 408.3 m/s F=ma = 11.2 a solve for a: \(V_f^2−V_o^2=2as\\(408.3)^2-0=2a(0.6)\\ \Large a=\frac{(408.3)^2}{2(0.6)}\) what did you get for a?

OpenStudy (jadzia):

@Joeyray still there?